### v1

# Bridging the Editing Gap in LLMs: FineEdit for Precise and Targeted Text Modifications

Abstract
Large Language Models (LLMs) have transformed natural language processing, yet they still struggle with direct text editing tasks that demand precise, context-aware modifications. While models like ChatGPT excel in text generation and analysis, their editing abilities often fall short, addressing only superficial issues rather than deeper structural or logical inconsistencies. In this work, we introduce a dual approach to enhance LLMs editing performance. First, we present InstrEditBench, a high-quality benchmark dataset comprising over 20,000 structured editing tasks spanning Wiki articles, LaTeX documents, code, and database Domain-specific Languages (DSL). InstrEditBench is generated using an innovative automated workflow that accurately identifies and evaluates targeted edits, ensuring that modifications adhere strictly to specified instructions without altering unrelated content. Second, we propose FineEdit, a specialized model trained on this curated benchmark. Experimental results demonstrate that FineEdit achieves significant improvements around 10% compared with Gemini on direct editing tasks, convincingly validating its effectiveness.
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, unlocking capabilities once thought unattainable. ChatGPT, for example, shows exceptional skills in text generation and logical reasoning [37]. Despite these impressive advancements, LLMs still face significant challenges in the underperformance of text editing tasks. [7] has mentioned that the use of ChatGPT for editing has obvious limitations, which might not accurately follow the editing task instructions and understand the author’s intent, leading to changes that are not appropriate to the context of the text. Meanwhile, it effectively addresses surface-level issues, such as spelling and formatting, but cannot resolve complex challenges, such as editing in long text context or strict following task instructions.
To address these limitations, researchers have developed methods to enhance the editing capabilities of LLMs, particularly under task-specific scenarios, e.g., editing in code, LaTeX, etc. However, LLMs’ general editing capabilities in task-specific settings often fall short [57, 31]. They tend to generate incorrect outputs and stray from the given editing instructions. This issue arises primarily because these models overly emphasize task-specific constraints and are susceptible to hallucinating extraneous information.
In contrast, we notice that if narrowing the model’s focus to just two factors: the precise location of the edit and the specific content to be changed, the edit task itself could be better accomplished. Per this intuition, we propose a dual approach consisting of a benchmark for editing tasks and a model named FineEdit. For the benchmark, we design an automated workflow that focuses on accurately identifying and evaluating structured text edits. This workflow identifies precise differences and ensures correct edits through quality control. By reducing noise and focusing on meaningful modifications, this process produces a benchmark that is both practical for training and robust for evaluation. It directly addresses limitations in existing methods and aligns better with the practical demands of real-world editing tasks.
Furthermore, we use part of the curated benchmark to train the model, focusing on direct editing tasks. FineEdit achieves an over 10% improvement over Gemini 1.5 Flash and Gemini 2.0 Flash [12], and up to 30% over Llama-3.2-3B [34] on diverse editing benchmarks, while outperforming Mistral-7B-OpenOrca [27, 35, 30] over 40% in direct editing tasks.
The main contributions of this work include:
  • A high-quality benchmark dataset (InstrEditBench) [^1] We will release all datasets and the code to promote reproducibility on acceptance.: We created a curated dataset with 20,000+ structured editing tasks across Wiki articles, LaTeX documents, Code, and Database DSL, providing a unified evaluation standard for structured text editing research.
  • An innovative automated dataset generation workflow: We developed a comprehensive workflow that ensures the benchmark’s quality by accurately identifying line numbers and applying rigorous criteria to filter meaningful and relevant edits.
  • The FineEdit model: We introduce a specialized model designed for structured direct text editing, demonstrating superior performance across benchmarks compared with existing models.
2 Background
2.1 Problem Formulation
Each data point consists of an original structured text, $T_{\text{orig}}$, and an editing instruction, $I_{\text{edit}}$. The objective is to generate an edited text, $T_{\text{edit}}$, that incorporates the modifications specified by $I_{\text{edit}}$. Formally, this process is defined as
$T_{\text{edit}}=f\Bigl{(}T_{\text{orig}},I_{\text{edit}};\theta\Bigr{)}$ (1)
where $\theta$ represents learned parameters and $f$ denotes a function instantiated by a LLM that maps the original text $T_{\text{orig}}$ and editing instruction $I_{\text{edit}}$ to the edited text $T_{\text{edit}}$.
The parameters $\theta$ are learned from a dataset consisting of triples $\{(T_{\text{orig}}^{(i)},I_{\text{edit}}^{(i)},T_{\text{edit}}^{(i)})\}_{i=1}^{N}$ during training, where the objective is to minimize the discrepancy between the generated output and the ground truth edited text.
Internally, $f$ concatenates $T_{\text{orig}}$ and $I_{\text{edit}}$ into a single prompt and generates $T_{\text{edit}}$ token by token in an autoregressive manner. Specifically, if $T_{\text{edit}}=(y_{1},y_{2},\dots,y_{t})$, the probability of the edited text is factorized as
$\begin{split}p(T_{\text{edit}}\mid T_{\text{orig}},I_{\text{edit}})=&\prod_{i=1}^{t}p\Bigl{(}y_{i}\mid T_{\text{orig}},I_{\text{edit}},\\ &y_{1},y_{2},\dots,y_{i-1}\Bigr{)}\end{split}$ (2)
For finetuning on the editing task, the prompt tokens (i.e., the original text and the editing instruction) are masked out in the loss function to ensure that the model focuses only on predicting the correct edited tokens. At inference time, the model processes the prompt and subsequently generates $T_{\text{edit}}$.
The parameters $\theta$ are fine-tuned on labeled examples $(T_{\text{orig}},I_{\text{edit}},T_{\text{edit}})$ by minimizing the negative log-likelihood of the target tokens with the loss:
$\mathcal{L}(\theta)=-\sum_{t=1}^{|T_{\text{edit}}|}\log P_{\theta}(y_{t}\mid T_{\text{orig}},I_{\text{edit}},y_{1:t-1})$ (3)
over all training samples in the dataset
2.2 LLM Editing Tasks
LLMs are increasingly recognized as versatile tools for automating and enhancing editing tasks across diverse domains. Previous studies have explored LLMs for editing tasks in areas such as natural language (e.g., wiki articles) and code. For instance, CoEdIT [39] employs task-specific instruction tuning to achieve precise modifications, while other works fine-tune models like T5 [38] on pairs of original and edited texts [16, 41, 32, 14, 15, 24]. However, many of these approaches rely on specialized techniques or focus narrowly on specific tasks, such as grammar correction [32, 18], text simplification [59], paraphrase generation [10], or style transfer [42], which limits their generalizability across a broader range of editing scenarios. In the realm of code editing, Fan et al. [17] examined LLMs for code change tasks and identified weaknesses in generating accurate reviews and commit messages. While these studies offer valuable insights, they often fall short in providing unified benchmarks and robust solutions to address the full spectrum of editing challenges. Our work addresses these gaps by introducing a comprehensive, cross-scenario editing tasks benchmark that covers Wiki, code, DSL, and LaTeX.
2.3 LLM Benchmarking
LLM benchmarking is a crucial aspect of evaluating the diverse capabilities of LLM. Researchers have developed numerous benchmarks spanning multiple domains, including code [8, 1, 23, 56], commonsense reasoning [3, 44, 58, 43], reading comprehension [40, 9, 11], and language understanding [47, 48, 54]. However, only a few works benchmark the editing performance of LLMs. For example, GEM [55] introduces metrics for subjective tasks without gold standards, while CriticBench [28] assesses iterative output refinement. Additionally, [5] explores that fine-tuning with curated training data significantly improves code editing performance. Some automated evaluation is also involved in LLM benchmarking. For instance, G-Eval [29] is an automated evaluation framework that leverages large language models to assess text quality and model performance in generative tasks. Built on a Chain-of-Thought (CoT) prompting strategy [51], G-Eval guides the model to articulate intermediate reasoning steps before reaching its final evaluation, leading to outputs that align closely with human judgments [29]. However, these efforts focus on short-context, isolated tasks and do not systematically evaluate an LLM’s ability to locate and modify content within long contexts. Our work addresses this gap by introducing a comprehensive benchmark covering Wiki, code, DSL, and LaTeX, emphasizing long-context editing.
Figure 1: Workflow of Generating High-quality FineEdit benchmark. The content difference is highlighted in red.
3 Method
3.1 Instruction categories
We leverage four data sources to cover a wide range of representative text application scenarios: Wiki, Code, DSL, and LaTeX. The details of each categories are described as follows:
  • Wiki: Data is extracted from the WikiText language modeling dataset [33], which contains over 100 million tokens from a dedicated subset of Wikipedia’s Good articles [53] and Wikipedia’s Featured articles [52]. Specifically, sections from these articles are extracted and then contiguous segments are randomly selected to provide data points with various lengths.
  • Code: Code samples are extracted from the CodeSearchNet corpus [22], which contains about two million pairs of comments and code from GitHub projects. To make the edit task more challenging, each code sample in our benchmark is made up of several instead of one code segment because one single code segment is too short (about 10 lines).
  • DSL: Database Domain Specific Language (DSL) is also considered in our benchmark. It consists of queries and schema definitions from multiple public repositories [20, 2, 4, 26].
  • LaTeX: LaTeX data is extracted from the Latex2Poster dataset [25] that offers the LaTeX source code document of research papers along with metadata. Specifically, each data point in our benchmark consists of multiple subsections from each extracted document data.
3.2 Instruction Generation
Zero-shot instruction generation is efficient but often lacks diversity. To address this limitation, we build on the work of [49, 46] by leveraging ChatGPT-4o mini combined with in-context learning (ICL) [13]. Our approach is designed to generate specific edit requests tailored to the structural characteristics of different data categories, as process ➀ in Figure 1. For Wiki, which primarily consists of clear structural text elements like headings and subheadings, we apply a zero-shot prompting strategy. In contrast, for more complex domains such as LaTeX, code, and DSL, we adopt ICL to improve the diversity and nuance of generated instructions. This category-specific strategy not only enriches the instruction sets but also enhances their ability to capture domain-specific editing challenges without compromising on precision and efficiency. We will describe prompt details in Appendix A.
3.3 Instruction filtering
After obtaining the edit instructions for each content, we apply them to the original text to produce an edited version as process ➁ in Figure 1. However, ensuring the quality of the edited content remains challenging. Although LLM generally follows the edit instructions, errors may occur—for example, targeting incorrect line numbers or misinterpreting the intended semantics [50, 6]. To address this problem and improve data quality, we propose DiffEval Pipeline, which integrates G-Eval [29] and Git-Diff as an automatic filter to improve data quality.
Besides adopting G-Eval for automated assessment [29], the DiffEval Pipeline also relies on git [19], a widely used version control system, to detect and classify textual modifications. Specifically, the command git diff specifies differences between the original and modified texts as process ➂ in Figure 1, categorizing changes into four types:
  • Replacements: an original segment is transformed into a new form, indicated as [original_text -> modified_text]. This captures cases where an existing text portion is substituted with different content, which may alter meaning or style.
  • Deletions: a segment is removed entirely, shown as [-original_text-]. Such removals can simplify the text or eliminate irrelevant or erroneous sections.
  • Insertions: new content is added, denoted as [+modified_text+]. Insertions enrich the text with extra details, clarifications, or elaborations.
  • Unchanged Text: labeled as equal: unchanged_text. This indicates portions that remain identical between the original and modified versions, providing a reference for what the model has chosen to retain.
By categorizing changes into these four types, the DiffEval Pipeline offers a structured view of how text is altered, enabling more precise evaluations when paired with G-Eval.
[…]
This indicates that the duplicate has been successfully removed.
Finally, process ➃ in Figure 1 demonstrates that DiffEval carefully reviews the aggregated data (marked with red arrows) alongside the edit request to fully grasp the context, structure, and nuances of the text. It identifies discrepancies between the intended edits and the actual modifications, verifying whether the changes faithfully implement the edit instructions. By using the git diff output instead of the complete edited content, DiffEval can precisely locate modifications using supplementary information such as line numbers and structured differences. Moreover, git diff minimizes unnecessary noise and reduces computational overhead by significantly lowering the token count compared with the full edited content. Once all required data is gathered, the G-Eval analysis process evaluates the collected information to further enhance the dataset quality.
Specifically, the analysis process begins by parsing the structure of git diff outputs, categorizing changes as replacements, deletions, insertions, or unchanged segments. Next, it evaluates the semantic meaning of both the original content and the modifications to ensure that the changes are accurate and complete. This involves a thorough review of the original text, the edit request, and the resulting edits, applying predefined categorization rules, and assessing overall coherence.
Based on this analysis process, the DiffEval is able to assign a coherence score, G-Score, to the edited content, reflecting the semantic integrity and logical consistency of the modifications. This score is used to filter out output that does not meet the desired quality threshold $\alpha$.
3.4 Data Statistics
Our curated benchmark comprises 28,050 items spanning a diverse array of structured data types, including 8,366 LaTeX contexts, 7,712 code segments, 8,025 WikiText entries, and 3,947 database language samples, thereby reflecting both the generality and scale of real-world structured data. Table 1 shows the example across four categories. For each item, it has the following attributes:
  • Id: a unique identifier for each entry.
  • Original content: the content directly extracted from the data source.
  • Edit request: The editing instruction generated through zero-shot or few-shot prompting based on the original content
  • Edited Content: the output after applying edit request to the original content.
  • Difference: the changed part between Edit content and original content.
  • G-score: evaluates the quality of the edited content based on its strict adherence to the edit request content.
Data Category Orignal Content Edit Request Edited Content Difference G-score
WikiText …As with previous <unk> Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a military unit… Replace “\<unk>\” with “Valkyria” where it appears in the text. …As with previous Valkyria Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a military unit… Line 2 differs: Differences: …As with previous [<un -> Val]k[> -> yria] Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a mili- tary unit… 9
[…]
Table 1: Data examples of different data categories with all attributes (content, edit request, edited content, difference, and G-score).
4 Evaluation
4.1 Experimental Setup
In this section, we detail the experimental setups, including dataset splits, model variants, baselines, evaluation metrics, and implementation specifics.
Dataset and Model Variants. We evaluate FineEdit on our proposed InstrEditBench using a 90/10 train-test split. Additionally, we introduce three versions of FineEdit—FineEdit-L, FineEdit-XL, and FineEdit-Pro—fine-tuned from LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen2.5-3B-Instruct base models, respectively, to cover a wide spectrum of architectures and parameter scales.
Baselines. Our baselines include Gemini 1.5 Flash, Gemini 2.0 Flash, LLaMA-3.2-1B, LLaMA-3.2-3B, Qwen2.5-3B-Instruct, and Mistral-7B, spanning diverse architectures and sizes. We evaluate both zero-shot and few-shot prompting on the Gemini models, while open-source models are assessed using zero-shot prompting.
Metrics. Following established approaches [36, 45], we use BLEU and ROUGE-L metrics to assess the vocabulary and structural consistency between the edited and reference texts.
Implementation details. For existing models, we strictly adhere to configurations from their original papers. To manage fixed maximum token lengths $L$, if the combined $T_{\text{orig}}$ and $I_{\text{edit}}$ exceed $L$, we partition $T_{\text{orig}}$ into chunks of size $\leq L$, process each chunk independently with the same edit instruction, and concatenate the outputs to form the complete edited text. We fine-tune models using Low-Rank Adaptation (LoRA) [21] with $r=8$, $\alpha=32$, and a dropout rate of 0.05, employing the AdamW optimizer with a learning rate of $2\times 10^{-5}$, training for 2 epochs, an effective batch size of 1, and 4 gradient accumulation steps. During generation, we set the temperature to 0.7 and use top-p sampling with $p=0.9$, merging outputs from all chunks to produce the final edited text. Additional hyperparameter configurations and training details are provided in Appendix B.
4.2 Performance of Existing Models
We evaluated FineEdit against several state-of-the-art baselines on the InstrEditBench dataset across four data categories as presented in Table 2.
Comparison with Zero-shot Performance. Among all baselines, Gemini 1.5 Flash achieved the highest overall scores, while Mistral-7B-OpenOrca recorded the lowest BLEU and ROUGE-L values. Although model size is often a crucial factor, Gemini 2.0 Flash did not surpass Gemini 1.5 Flash in overall effectiveness. For instance, despite having more parameters than LLaMA-3.2-1B, Mistral-7B-OpenOrca underperformed in both metrics, highlighting the significance of model architecture and training methods. Moreover, while Gemini 2.0 Flash shows superior semantic understanding in the Wiki category—achieving a BLEU score of 0.9133 and a ROUGE-L score of 0.9429—its overall performance remains below that of its counterpart.
FineEdit, and in particular its FineEdit-Pro variant, further outperforms all zero-shot baselines. FineEdit-Pro achieves an overall BLEU score of 0.9245, representing improvements of approximately 11.6%, 57.7%, and 184.7% over Gemini 1.5 Flash (0.8285), LLaMA-3.2-3B (0.5862), and Mistral-7B-OpenOrca (0.3246), respectively. These gains are consistently observed across individual data categories—for example, FineEdit-Pro attains BLEU scores of 0.9521 and 0.9538 in the DSL and Code domains, respectively. These results underscore the effectiveness of FineEdit’s targeted fine-tuning strategy, which focuses on precise editing of location and content to preserve both structural and semantic integrity.
Comparison with Few-shot Performance. We further evaluated few-shot learning on the Gemini models. Although few-shot performance notably improved in some categories—for example, in the LaTeX domain, where Gemini 2.0 Flash exhibited a 20% higher BLEU score than in the zero-shot setting—the overall few-shot results still lag behind FineEdit. In certain cases, such as the SQL category, few-shot learning made little difference, with BLEU and ROUGE-L scores of only 0.1600 and 0.1814, respectively. These findings reinforce the value of our curated benchmark in driving improvements in editing tasks.
Key Findings: FindEdit demonstrates robust overall effectiveness across Wiki, Code, DSL, and LaTeX categories. These results not only position FineEdit as a competitive method for structured editing tasks but also provide valuable insights into how targeted training strategies can elevate model performance in diverse application scenarios.
Method Model Open-Source LaTeX DSL Wiki Code Overall BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L Zero-shot Gemini 1.5 Flash 0.8665 0.9150 0.8297 0.8555 0.7626 0.8361 0.8551 0.9073 0.8285 0.8819 Gemini 2.0 Flash 0.7413 0.7951 0.4706 0.4964 0.9133 0.9429 0.1339 0.2737 0.5853 0.6519 Llama-3.2-1B 0.5088 0.6108 0.5564 0.6596 0.4413 0.5766 0.4742 0.6072 0.4867 0.6069 Llama-3.2-3B 0.5969 0.6925 0.5747 0.6821 0.5061 0.6384 0.6638 0.7727 0.5862 0.6976 Qwen2.5-3B-Instruct 0.5467 0.6712 0.4107 0.4991 0.4170 0.5699 0.3967 0.5390 0.4492 0.5816 Mistral-7B-OpenOrca 0.3782 0.5770 0.0361 0.1638 0.3608 0.5840 0.3763 0.6447 0.3246 0.5395 Few-shot Gemini 1.5 Flash(2-shot) 0.8742 0.9324 0.0908 0.1190 0.8657 0.9139 0.7412 0.8302 0.7249 0.7845 Gemini 2.0 Flash(2-shot) 0.9464 0.9723 0.1600 0.1814 0.9380 0.9665 0.8327 0.8698 0.8011 0.8302 FineEdit FineEdit-L 0.9311 0.9697 0.9334 0.9615 0.8077 0.9036 0.9296 0.9725 0.8957 0.9504 FineEdit-XL 0.8867 0.9502 0.9241 0.9552 0.8120 0.9056 0.9295 0.9720 0.8824 0.9441 FineEdit-Pro 0.9539 0.9821 0.9521 0.9710 0.8521 0.9185 0.9538 0.9836 0.9245 0.9628
Table 2: Comparison of LLMs on BLEU and ROUGE-L for LaTeX, DSL, Wiki, Code. Overall data displays average performance among all data categories. The best results are highlighted in bold.
4.3 FineEdit: Supervised Finetuning
Our FineEdit model is offered in three variants: FineEdit-L, FineEdit-XL, and FineEdit-Pro. Under zero-shot conditions, FineEdit-L consistently outperforms all baseline models in BLEU and ROUGE-L scores for LaTeX, DSL, Wiki, and Code tasks. For example, compared to Gemini 1.5 Flash, FineEdit-L improves overall BLEU scores by roughly 8%, with even larger gains observed in specific categories. Notably, FineEdit-XL performs similarly to FineEdit-L, suggesting that increasing the parameter count from 1B to 3B using LLaMA does not yield a significant performance boost.
By leveraging the superior instruction-following capabilities of Qwen2.5-3B-Instruct, our final variant, FineEdit-Pro, further elevates performance. FineEdit-Pro achieves an overall BLEU score of 0.9245, which represents improvements of approximately 11.6% over Gemini 1.5 Flash, and gains of around 14.7% and 11.7% in the DSL and Wiki tasks, respectively. These consistent improvements across multiple data categories underscore the effectiveness of our supervised fine-tuning strategy and highlight the importance of a strong instruction-tuned base model over merely increasing model size.
We also compared our models with Gemini’s few-shot prompting approach in real-world scenarios. Although in-context learning (ICL) boosts Gemini’s performance in some cases—such as a 8% higher BLEU score in Wiki dataset for Gemini 2.0 Flash—the overall results still lag behind FineEdit-Pro. This evidence confirms that our tailored supervised fine-tuning approach yields a more robust and generalizable solution for structured editing tasks.
Key Findings: FineEdit’s supervised fine-tuning markedly enhances performance. FineEdit-L surpasses zero-shot baselines and FineEdit-XL offers comparable gains, while FineEdit-Pro (built on Qwen2.5-3B-Instruct) achieves the highest scores. This highlights that robust instruction tuning is more effective than merely scaling model size.
4.4 Qualitative Study
To qualitatively assess the performance of FindEdit, we conduct several studies as shown in Figure 2. This figure illustrates eight examples of how FineEdit-Pro and Gemini respond to diverse editing requests. In several cases, FineEdit-Pro accurately applies changes—such as adding new columns in DSL or adjusting environment commands—while Gemini often restates the instruction without implementing the intended modifications.
Specifically, both Gemini 1.5 Flash and 2.0 Flash perform well on LaTeX and Wiki tasks, yet they struggle with DSL and Code tasks. For example, as shown in Figure 2, FineEdit-Pro correctly identifies the target table and appends a new column named created_at with the data type DEFAULT CURRENT_TIMESTAMP. In contrast, Gemini misinterprets the instruction, merely repeating the edit request rather than applying the intended change. These observations highlight the qualitative strengths of our proposed FineEdit approach.
Nonetheless, FineEdit is not without shortcomings. In the LaTeX example depicted in Figure 2, Gemini accurately locates the subsection{Strengths} and updates it as specified. However, although FineEdit-Pro also identifies and modifies the correct location, it generates the correct response twice, which deviates from the direct editing requirement. This discrepancy suggests that FineEdit-Pro, though generally more reliable, can overapply modifications in specific cases.
Overall, these results illustrate FineEdit-Pro’s capacity to handle more complex edits, particularly for DSL and Code, while Gemini often fails to implement them. Nevertheless, occasional issues like duplicate outputs highlight the need for refinement, ensuring FineEdit-Pro consistently adheres to direct editing requirements without introducing redundant content. On the other hand, Gemini occasionally performs better in simpler tasks, such as LaTeX updates.
Key Findings: FineEdit-Pro demonstrates superior handling of DSL and Code edits compared to Gemini, though minor issues such as duplicate outputs in LaTeX tasks remain. Overall, FineEdit’s qualitative performance confirms its robust ability to interpret and execute complex editing instructions.
Figure 2: Comparison between Gemini and FindEdit Pro response.
4.5 Human Evaluation
To assess whether DiffEval enhances overall dataset quality, we conducted a human evaluation. Given that our dataset includes Code and DSL categories—areas closely tied to computer programming—we have three evaluators, each holding at least a Bachelor’s degree in a Computer Science-related field. We established the following guidelines to ensure rigorous assessment: (1) Precise Observation: Confirm that the updated content exactly corresponds to the segment specified by the edit request. (2) No Unintended Modifications: Verify that no other sections have been altered; any unexpected changes result in failure. (3) Three-Round Procedure: Two evaluators independently review each item, with a third evaluator resolving any discrepancies.
We examined 100 items per category and found that data processed through our DiffEval pipeline exhibited noticeably enhanced accuracy, as shown in Table 3. The Wiki and Code datasets, in particular, demonstrated the most reliable outcomes, with edited content precisely matching the requested modifications. Notably, the DSL dataset experienced the greatest improvement, with quality increasing by over 24% compared to data that did not meet DiffEval’s standards.
Wiki LaTeX DSL Code
G-score $\geq$ 9 97% 93% 90% 97%
G-score $<$ 9 87% 89% 66% 83%
Table 3: Sample performance based on the G-Score
Key Findings: The DiffEval pipeline significantly improves dataset quality, with Wiki and Code categories achieving high precision, and the DSL category showing over 24% enhancement in quality.
5 Conclusion
This work addresses the critical gap in LLMs’ ability to perform precise and targeted text modifications. We introduce InstrEditBench, a high-quality benchmark with 20,000+ structured editing tasks across Wiki articles, LaTeX documents, code, and database DSLs, enabling rigorous evaluation of direct editing capabilities. To further advance LLMs’ editing proficiency, we propose FineEdit, a specialized model trained on this benchmark. Extensive evaluations demonstrate that FineEdit outperforms state-of-the-art models, including GPT-4o, Gemini 2.0, and LLaMa-3.2, with up to 10% improvement compared to Gemini in direct editing task performance.
6 Limitations
Limited Deployment Scope. Due to cost and hardware constraints, our evaluations were limited to large proprietary LLMs (e.g., Gemini), rather than large open-source models.
Controlled Context Evaluation. Our benchmark focuses on controlled evaluation context, where it does not yet encompass long-context chain-of-thought scenarios, as smaller LLMs are confined by limited context windows, even though such techniques could be effective in proprietary models
Appendix A Dataset Generation Prompts
We use the following prompts for dataset generation on each domain.
user_prompt = r’’’Task: Gener ate one precise editing request for the given LaTeX code, focusing exclusively on one detailed LaTeX-specific aspect.
     1. Analyze LaTeX Components: Examine the LaTeX code thoroughly, identifying elements such as commands, environments, packages, mathematical expressions, figures, tables, references, labels, and syntax structures.
     2. Target a Single LaTeX Issue: The editing request must address only one specific LaTeX-related issue such as commands, environments, packages, mathematical expressions, figures, tables, references, labels, and syntax structures.
     3. Clearly define the exact edit needed. The action should be definitive and unambiguous, avoiding any form of suggestion, optional language, or choices. Do not include reasons for the edit or any additional information beyond the request.
     4. Do not include reasons for the edit or any additional information beyond the edit request. The request should be a direct instruction.
     The request examples are:
     [Example 1]
     <Edit Request>
[…]
     </Edit Request>
     [Example 2]
     <Edit Request>
[…]
     </Edit Request>
     [Example 3]
     <Edit Request>
[…]
     </Edit Request>
     [Example 4]
     <Edit Request>
     Change the column specification in the tabular environment from {l l l} to {l c r} to adjust the alignment of the data columns.
     </Edit Request>
     [Example 5]
     <Edit Request>
[…]
     </Edit Request>
     [Example 6]
     <Edit Request>
[…]
     </Edit Request>
[Example 7]
     <Edit Request>
[…]
     {example-image} to scale the image.
     </Edit Request>
     [Example 8]
     <Edit Request>
[…]
     </Edit Request>
     I will give you the content and then the editing request.
     Please Edit the content based on the editing request.
     While Editing, don’t add other words like
     modified or something. Just Edit directly.
     Content: {original_context}
     Editing Request: {edit_request}
     Please return the complete content after editing.
     Don’t skip the empty line and keep the original
     apart from the editing part.
We use the following prompt for doing G-eval.
Appendix B Additional Implementation Details
Chunking long context: Many large language models impose a fixed maximum token length $L$ on their input (and sometimes output) sequences. Consequently, if the combination of $T_{\text{orig}}$ and $I_{\text{edit}}$ exceeds this limit, we divide the $T_{\text{orig}}$ into smaller chunks of size $\leq L$. Each chunk is then processed independently—paired with the same edit request and later concatenated to form the complete edited text. This approach ensures that every chunk fits within the model’s token budget, preventing overflow and reducing memory usage while preserving the overall structured editing behavior.
Fine-Tuning Strategy: We use Low-Rank Adaptation (LoRA) [21] to efficiently adapt these models to our task, significantly reducing the number of trainable parameters while preserving their expressive power. In all LoRA configurations, We set the rank $r=8$ and scaling $\alpha=32$, and use a dropout probability of 0.05. For both Llama-based and Qwen-based models, we apply LoRA to the attention’s projection layers through trainable low-rank matrices. We used the AdamW optimizer with a learning rate of $2\times 10^{-5}$, training for 2 epochs, and set the effective batch size of 1 with gradient accumulation steps of 4 due to device limits. This strategy not only reduces computational overhead but also enables rapid convergence on our structured editing tasks. Preliminary experiments guided the choice of hyperparameters across all three model variants.
Decoding and Inference: During generation, we set the temperature to 0.7 and used top-p sampling with a probability of 0.9 to balance diversity and coherence. Greedy decoding is applied by default if without sampling setting. The final edited text is obtained by merging the edited outputs from all chunks.

### v2

@@
-Bridging the Editing Gap in LLMs: FineEdit for Precise and Targeted Text Modifications
+FineEdit: Unlock Instruction-Based Text Editing for LLMs
 Abstract
-Large Language Models (LLMs) have transformed natural language processing, yet they still struggle with direct text editing tasks that demand precise, context-aware modifications.
-While models like ChatGPT excel in text generation and analysis, their editing abilities often fall short, addressing only superficial issues rather than deeper structural or logical inconsistencies.
-In this work, we introduce a dual approach to enhance LLMs editing performance.
-First, we present InstrEditBench, a high-quality benchmark dataset comprising over 20,000 structured editing tasks spanning Wiki articles, LaTeX documents, code, and database Domain-specific Languages (DSL).
-InstrEditBench is generated using an innovative automated workflow that accurately identifies and evaluates targeted edits, ensuring that modifications adhere strictly to specified instructions without altering unrelated content.
-Second, we propose FineEdit, a specialized model trained on this curated benchmark.
-Experimental results demonstrate that FineEdit achieves significant improvements around 10% compared with Gemini on direct editing tasks, convincingly validating its effectiveness.
+Large Language Models (LLMs) have significantly advanced natural language processing, demonstrating strong capabilities in tasks such as text generation, summarization, and reasoning.
+Recently, their potential for automating precise text editing tasks across specialized domains, such as programming code, LaTeX, and structured database languages, has gained attention.
+However, current state-of-the-art LLMs still struggle with executing precise, instruction-driven edits, particularly when structural accuracy and strict adherence to domain conventions are required.
+To address these challenges, we introduce InstrEditBench, an automated benchmark dataset comprising over 30,000 structured editing tasks spanning diverse domains, including Wikipedia articles, LaTeX documents, source code, and database languages.
+Using this benchmark, we develop FineEdit, a specialized editing model explicitly trained for accurate, context-aware text modifications.
+Experimental evaluations demonstrate that FineEdit outperforms state-of-the-art models, achieving improvements of approximately 10% over Gemini models on single-turn edits, up to 30% over Llama-3.2-3B, and exceeding Mistral-7B-OpenOrca performance by over 40% on direct editing tasks.
+FineEdit also effectively generalizes to realistic multi-turn editing scenarios, highlighting its practical applicability.
+FineEdit: Unlock Instruction-Based Text Editing for LLMs
+Yiming Zeng1 Wanhao Yu2 Zexin Li3 Tao Ren4 Yu Ma5 Jinghan Cao6 Xiyan Chen4 Tingting Yu1 1University of Connecticut 2University of North Carolina at Charlotte 3University of California, Riverside 4University of Pittsburgh 5Carnegie Mellon University 6San Francisco State University
 1 Introduction
-Large Language Models (LLMs) have revolutionized natural language processing, unlocking capabilities once thought unattainable.
-ChatGPT, for example, shows exceptional skills in text generation and logical reasoning [37].
-Despite these impressive advancements, LLMs still face significant challenges in the underperformance of text editing tasks.
-[7] has mentioned that the use of ChatGPT for editing has obvious limitations, which might not accurately follow the editing task instructions and understand the author’s intent, leading to changes that are not appropriate to the context of the text.
-Meanwhile, it effectively addresses surface-level issues, such as spelling and formatting, but cannot resolve complex challenges, such as editing in long text context or strict following task instructions.
-To address these limitations, researchers have developed methods to enhance the editing capabilities of LLMs, particularly under task-specific scenarios, e.g., editing in code, LaTeX, etc.
-However, LLMs’ general editing capabilities in task-specific settings often fall short [57, 31].
+Large Language Models (LLMs) have brought transformative progress to the field of natural language processing, demonstrating remarkable capabilities in text generation, summarization, and reasoning Brown et al.
+(2020); Achiam et al.
+(2023).
+Recently, LLMs have received increasing attention for their potential to automate and enhance text editing across a variety of domains Celikyilmaz et al.
+(2020).
+Such editing capabilities is particularly needed under task-specific application scenarios, e.g., code editing Fan et al.
+(2024), Wiki editing Suri et al.
+(2024), etc.
+Despite this promise, current LLMs still face notable limitations when applied to tasks that demand direct editing, where the model must simultaneously understand the original text, follow the instruction precisely, and generate semantically aligned, high-quality edits.
+Even powerful proprietary tools like ChatGPT often struggle to fully understand user intent and reliably follow strict editing instructions, especially in long-context scenarios Castillo-González et al.
+(2022).
+Particularly, LLMs’ general editing capabilities in task-specific settings often fall short Yao et al.
+(2023); Ma et al.
+(2024).
 They tend to generate incorrect outputs and stray from the given editing instructions.
-This issue arises primarily because these models overly emphasize task-specific constraints and are susceptible to hallucinating extraneous information.
-In contrast, we notice that if narrowing the model’s focus to just two factors: the precise location of the edit and the specific content to be changed, the edit task itself could be better accomplished.
-Per this intuition, we propose a dual approach consisting of a benchmark for editing tasks and a model named FineEdit.
-For the benchmark, we design an automated workflow that focuses on accurately identifying and evaluating structured text edits.
+To address these challenges, we propose a more focused approach to editing with LLMs.
+Our key insight is that narrowing the model’s attention to two fundamental aspects, the exact location of the edit and the content to be modified, can significantly improve performance in direct editing tasks.
+Per this intuition, we propose a dual approach consisting of a dedicated benchmark (InstrEditBench) for editing tasks and an editing-specific model (FineEdit).
+Specifically, we design an automated workflow that focuses on accurately identifying and evaluating structured text edits.
 This workflow identifies precise differences and ensures correct edits through quality control.
-By reducing noise and focusing on meaningful modifications, this process produces a benchmark that is both practical for training and robust for evaluation.
+By reducing noise and focusing on meaningful modifications, this process produces a dedicated, high-quality benchmark.
 It directly addresses limitations in existing methods and aligns better with the practical demands of real-world editing tasks.
-Furthermore, we use part of the curated benchmark to train the model, focusing on direct editing tasks.
-FineEdit achieves an over 10% improvement over Gemini 1.5 Flash and Gemini 2.0 Flash [12], and up to 30% over Llama-3.2-3B [34] on diverse editing benchmarks, while outperforming Mistral-7B-OpenOrca [27, 35, 30] over 40% in direct editing tasks.
+Notably, our approach is also generalized to multi-turn editing scenarios, a much more realistic user scenario, where instructions arrive iteratively and allow the model to refine its edits step by step.
+Implementation and evaluation.
+We train the FineEdit model on InstrEditBench benchmark, explicitly designed to optimize performance on direct, instruction-driven text editing tasks.
+The result shows that FineEdit achieves an improvement of 10% over Gemini 1.5 Flash and Gemini 2.0 Flash DeepMind (2024) in single-turn editing tasks, and up to 30% over Llama-3.2-3B Meta AI (2024) on diverse editing benchmarks, while outperforming Mistral-7B-OpenOrca Lian et al.
+(2023); Mukherjee et al.
+(2023); Longpre et al.
+(2023) over 40% on direct editing tasks.
 The main contributions of this work include:
-  • A high-quality benchmark dataset (InstrEditBench) [^1] We will release all datasets and the code to promote reproducibility on acceptance.: We created a curated dataset with 20,000+ structured editing tasks across Wiki articles, LaTeX documents, Code, and Database DSL, providing a unified evaluation standard for structured text editing research.
-  • An innovative automated dataset generation workflow: We developed a comprehensive workflow that ensures the benchmark’s quality by accurately identifying line numbers and applying rigorous criteria to filter meaningful and relevant edits.
-  • The FineEdit model: We introduce a specialized model designed for structured direct text editing, demonstrating superior performance across benchmarks compared with existing models.
+  • A high-quality benchmark (InstrEditBench) [^1] We will release all datasets and the code to promote reproducibility on acceptance.: We introduce the first systematically constructed benchmark that spans four diverse domains and contains more than 30,000 single-turn and multi-turn structured editing tasks, thereby establishing a unified and comprehensive evaluation standard for direct editing research.
+  • An innovative automated dataset generation workflow: We develop a comprehensive workflow that ensures the benchmark’s quality by accurately identifying line numbers and applying rigorous criteria to filter meaningful and relevant edits.
+  • The FineEdit model: We present a specialized model designed for direct text editing, demonstrating superior performance across benchmarks compared with existing models.
 2 Background
@@
 $\mathcal{L}(\theta)=-\sum_{t=1}^{|T_{\text{edit}}|}\log P_{\theta}(y_{t}\mid T_{\text{orig}},I_{\text{edit}},y_{1:t-1})$ (3)
-over all training samples in the dataset
+over all training samples in the dataset.
 2.2 LLM Editing Tasks
@@
 Previous studies have explored LLMs for editing tasks in areas such as natural language (e.g., wiki articles) and code.
-For instance, CoEdIT [39] employs task-specific instruction tuning to achieve precise modifications, while other works fine-tune models like T5 [38] on pairs of original and edited texts [16, 41, 32, 14, 15, 24].
-However, many of these approaches rely on specialized techniques or focus narrowly on specific tasks, such as grammar correction [32, 18], text simplification [59], paraphrase generation [10], or style transfer [42], which limits their generalizability across a broader range of editing scenarios.
+For instance, CoEdIT Raheja et al.
+(2023) employs task-specific instruction tuning to achieve precise modifications, while other works fine-tune models like T5 Raffel et al.
+(2020) on pairs of original and edited texts Faltings et al.
+(2021); Reid and Neubig (2022); Mallinson et al.
+(2022); Du et al.
+(2022a, b); Kim et al.
+(2022).
+However, many of these approaches rely on specialized techniques or focus narrowly on specific tasks, such as grammar correction Mallinson et al.
+(2022); Fang et al.
+(2023), text simplification Štajner et al.
+(2022), paraphrase generation Chowdhury et al.
+(2022), or style transfer Reif et al.
+(2022), which limits their generalizability across a broader range of editing scenarios.
 In the realm of code editing, Fan et al.
-[17] examined LLMs for code change tasks and identified weaknesses in generating accurate reviews and commit messages.
+Fan et al.
+(2024) examined LLMs for code change tasks and identified weaknesses in generating accurate reviews and commit messages.
+Beyond single-turn editing, iterative or multi-turn editing can further improve output quality by allowing incorporation of progressive feedback, leading to more accurate and context-aligned modifications Madaan et al.
+(2023); Schick et al.
+(2022).
 While these studies offer valuable insights, they often fall short in providing unified benchmarks and robust solutions to address the full spectrum of editing challenges.
 Our work addresses these gaps by introducing a comprehensive, cross-scenario editing tasks benchmark that covers Wiki, code, DSL, and LaTeX.
-2.3 LLM Benchmarking
-LLM benchmarking is a crucial aspect of evaluating the diverse capabilities of LLM.
-Researchers have developed numerous benchmarks spanning multiple domains, including code [8, 1, 23, 56], commonsense reasoning [3, 44, 58, 43], reading comprehension [40, 9, 11], and language understanding [47, 48, 54].
-However, only a few works benchmark the editing performance of LLMs.
-For example, GEM [55] introduces metrics for subjective tasks without gold standards, while CriticBench [28] assesses iterative output refinement.
-Additionally, [5] explores that fine-tuning with curated training data significantly improves code editing performance.
-Some automated evaluation is also involved in LLM benchmarking.
-For instance, G-Eval [29] is an automated evaluation framework that leverages large language models to assess text quality and model performance in generative tasks.
-Built on a Chain-of-Thought (CoT) prompting strategy [51], G-Eval guides the model to articulate intermediate reasoning steps before reaching its final evaluation, leading to outputs that align closely with human judgments [29].
-However, these efforts focus on short-context, isolated tasks and do not systematically evaluate an LLM’s ability to locate and modify content within long contexts.
-Our work addresses this gap by introducing a comprehensive benchmark covering Wiki, code, DSL, and LaTeX, emphasizing long-context editing.
-Figure 1: Workflow of Generating High-quality FineEdit benchmark. The content difference is highlighted in red.
+Figure 1: Workflow of Generating High-quality InstrEditBench. The content difference is highlighted in blue.
 3 Method
@@
 We leverage four data sources to cover a wide range of representative text application scenarios: Wiki, Code, DSL, and LaTeX. The details of each categories are described as follows:
-  • Wiki: Data is extracted from the WikiText language modeling dataset [33], which contains over 100 million tokens from a dedicated subset of Wikipedia’s Good articles [53] and Wikipedia’s Featured articles [52].
+  • Wiki: Data is extracted from the WikiText language modeling dataset Merity et al.
+(2016), which contains over 100 million tokens from a dedicated subset of Wikipedia’s Good articles Wikipedia (n.d.b) and Wikipedia’s Featured articles Wikipedia (n.d.a).
 Specifically, sections from these articles are extracted and then contiguous segments are randomly selected to provide data points with various lengths.
-  • Code: Code samples are extracted from the CodeSearchNet corpus [22], which contains about two million pairs of comments and code from GitHub projects.
+  • Code: Code samples are extracted from the CodeSearchNet corpus Husain et al.
+(2019), which contains about two million pairs of comments and code from GitHub projects.
 To make the edit task more challenging, each code sample in our benchmark is made up of several instead of one code segment because one single code segment is too short (about 10 lines).
-  • DSL: Database Domain Specific Language (DSL) is also considered in our benchmark. It consists of queries and schema definitions from multiple public repositories [20, 2, 4, 26].
-  • LaTeX: LaTeX data is extracted from the Latex2Poster dataset [25] that offers the LaTeX source code document of research papers along with metadata.
+  • DSL: Database Domain Specific Language (DSL) is also considered in our benchmark.
+It consists of queries and schema definitions from multiple public repositories hive (2024); b mc2 (2023); cassandra (2024); Lerocha (2024).
+  • LaTeX: LaTeX data is extracted from the Latex2Poster dataset Latex2Poster (2024) that offers the LaTeX source code document of research papers along with metadata.
 Specifically, each data point in our benchmark consists of multiple subsections from each extracted document data.
 3.2 Instruction Generation
-Zero-shot instruction generation is efficient but often lacks diversity.
-To address this limitation, we build on the work of [49, 46] by leveraging ChatGPT-4o mini combined with in-context learning (ICL) [13].
+Instruction Generation
+Zero-shot instruction generation is efficient, but often lacks diversity.
+To address this limitation, we build on the work of Wang et al.
+(2022); Taori et al.
+(2023) by leveraging ChatGPT-4o mini combined with in-context learning (ICL) Dong et al.
+(2024).
 Our approach is designed to generate specific edit requests tailored to the structural characteristics of different data categories, as process ➀ in Figure 1.
@@
 This category-specific strategy not only enriches the instruction sets but also enhances their ability to capture domain-specific editing challenges without compromising on precision and efficiency.
-We will describe prompt details in Appendix A.
+We will describe prompt details in Appendix C.
 3.3 Instruction filtering
@@
 However, ensuring the quality of the edited content remains challenging.
-Although LLM generally follows the edit instructions, errors may occur—for example, targeting incorrect line numbers or misinterpreting the intended semantics [50, 6].
-To address this problem and improve data quality, we propose DiffEval Pipeline, which integrates G-Eval [29] and Git-Diff as an automatic filter to improve data quality.
-Besides adopting G-Eval for automated assessment [29], the DiffEval Pipeline also relies on git [19], a widely used version control system, to detect and classify textual modifications.
+Although LLM generally follows the edit instructions, errors may occur—for example, targeting incorrect line numbers or misinterpreting the intended semantics Wang et al.
+(2025); Cassano et al.
+(2024).
+To address this problem and improve data quality, we propose DiffEval Pipeline, which integrates G-Eval Liu et al.
+(2023) and Git-Diff as an automatic filter to improve data quality.
+Besides adopting G-Eval for automated assessment (Liu et al., 2023), the DiffEval Pipeline also relies on git (git, 2024), a widely used version control system, to detect and classify textual modifications.
 Specifically, the command git diff specifies differences between the original and modified texts as process ➂ in Figure 1, categorizing changes into four types:
@@
 By categorizing changes into these four types, the DiffEval Pipeline offers a structured view of how text is altered, enabling more precise evaluations when paired with G-Eval.
-[…]
-[…]
-This indicates that the duplicate has been successfully removed.
 Finally, process ➃ in Figure 1 demonstrates that DiffEval carefully reviews the aggregated data (marked with red arrows) alongside the edit request to fully grasp the context, structure, and nuances of the text.
@@
 This involves a thorough review of the original text, the edit request, and the resulting edits, applying predefined categorization rules, and assessing overall coherence.
-Based on this analysis process, the DiffEval is able to assign a coherence score, G-Score, to the edited content, reflecting the semantic integrity and logical consistency of the modifications.
+Based on this analysis process, the DiffEval can assign a coherence score, G-Score, to the edited content, reflecting the semantic integrity and logical consistency of the modifications.
 This score is used to filter out output that does not meet the desired quality threshold $\alpha$.
-3.4 Data Statistics
-Our curated benchmark comprises 28,050 items spanning a diverse array of structured data types, including 8,366 LaTeX contexts, 7,712 code segments, 8,025 WikiText entries, and 3,947 database language samples, thereby reflecting both the generality and scale of real-world structured data.
-Table 1 shows the example across four categories.
-For each item, it has the following attributes:
-  • Id: a unique identifier for each entry.
-  • Original content: the content directly extracted from the data source.
-  • Edit request: The editing instruction generated through zero-shot or few-shot prompting based on the original content
-  • Edited Content: the output after applying edit request to the original content.
-  • Difference: the changed part between Edit content and original content.
-  • G-score: evaluates the quality of the edited content based on its strict adherence to the edit request content.
-Data Category Orignal Content Edit Request Edited Content Difference G-score
-WikiText …As with previous <unk> Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a military unit… Replace “\<unk>\” with “Valkyria” where it appears in the text.
-…As with previous Valkyria Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a military unit… Line 2 differs: Differences: …As with previous [<un -> Val]k[> -> yria] Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a mili- tary unit… 9
-[…]
-[…]
-[…]
-Table 1: Data examples of different data categories with all attributes (content, edit request, edited content, difference, and G-score).
-4 Evaluation
+3.4 Generalize to Multi-turn Editing Task
+Notably, our proposed framework is easily generalized to more practical multi-turn editing scenarios.
+Specifically, given the initial content, we instruct ChatGPT to generate a sequence of multiple distinct editing requests that are explicitly constrained to be non-contradictory with each other.
+Each generated editing request targets different aspects or details within the same content, ensuring that subsequent instructions complement rather than conflict with previous edits.
+This setting reflects real-world editing workflows where users iteratively refine content through consecutive instructions.
+4 Experiment
 4.1 Experimental Setup
@@
 Dataset and Model Variants.
-We evaluate FineEdit on our proposed InstrEditBench using a 90/10 train-test split.
-Additionally, we introduce three versions of FineEdit—FineEdit-L, FineEdit-XL, and FineEdit-Pro—fine-tuned from LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen2.5-3B-Instruct base models, respectively, to cover a wide spectrum of architectures and parameter scales.
+We evaluate FineEdit on our proposed InstrEditBench dataset using a 90/10 train-test split.
+Additionally, we introduce three FineEdit variants, namely FineEdit-L, FineEdit-XL, and FineEdit-Pro, which are fine-tuned from the LLaMA-3.2-1B, LLaMA-3.2-3B, and Qwen2.5-3B-Instruct base models respectively, covering a broad spectrum of model architectures and parameter scales.
 Baselines.
@@
 We evaluate both zero-shot and few-shot prompting on the Gemini models, while open-source models are assessed using zero-shot prompting.
-Metrics. Following established approaches [36, 45], we use BLEU and ROUGE-L metrics to assess the vocabulary and structural consistency between the edited and reference texts.
-Implementation details.
-For existing models, we strictly adhere to configurations from their original papers.
-To manage fixed maximum token lengths $L$, if the combined $T_{\text{orig}}$ and $I_{\text{edit}}$ exceed $L$, we partition $T_{\text{orig}}$ into chunks of size $\leq L$, process each chunk independently with the same edit instruction, and concatenate the outputs to form the complete edited text.
-We fine-tune models using Low-Rank Adaptation (LoRA) [21] with $r=8$, $\alpha=32$, and a dropout rate of 0.05, employing the AdamW optimizer with a learning rate of $2\times 10^{-5}$, training for 2 epochs, an effective batch size of 1, and 4 gradient accumulation steps.
-During generation, we set the temperature to 0.7 and use top-p sampling with $p=0.9$, merging outputs from all chunks to produce the final edited text.
-Additional hyperparameter configurations and training details are provided in Appendix B.
+Metrics.
+Following established approaches Nakamachi et al.
+(2020); Shen et al.
+(2017), we use BLEU and ROUGE-L metrics to assess the vocabulary and structural consistency between the edited and reference texts.
+Implementation details. Training details are provided in Appendix A.
+Method Model Open-Source LaTeX DSL Wiki Code Overall BLEU RG-L BLEU RG-L BLEU RG-L BLEU RG-L BLEU RG-L Zero-shot Gemini 1.5 Flash 0.8665 0.9150 0.8297 0.8555 0.7626 0.8361 0.8551 0.9073 0.8285 0.8819 Gemini 2.0 Flash 0.7413 0.7951 0.4706 0.4964 0.9133 0.9429 0.1339 0.2737 0.5853 0.6519 Llama-3.2-1B 0.5088 0.6108 0.5564 0.6596 0.4413 0.5766 0.4742 0.6072 0.4867 0.6069 Llama-3.2-3B 0.5969 0.6925 0.5747 0.6821 0.5061 0.6384 0.6638 0.7727 0.5862 0.6976 Qwen2.5-3B-Instr 0.5467 0.6712 0.4107 0.4991 0.4170 0.5699 0.3967 0.5390 0.4492 0.5816 Mistral-7B-Orca 0.3782 0.5770 0.0361 0.1638 0.3608 0.5840 0.3763 0.6447 0.3246 0.5395 Few-shot Gemini 1.5 Flash (2 shot) 0.8742 0.9324 0.0908 0.1190 0.8657 0.9139 0.7412 0.8302 0.7249 0.7845 Gemini 2.0 Flash (2 shot) 0.9464 0.9723 0.1600 0.1814 0.9380 0.9665 0.8327 0.8698 0.8011 0.8302 FineEdit FineEdit-L 0.9311 0.9697 0.9334 0.9615 0.8077 0.9036 0.9296 0.9725 0.8957 0.9504 FineEdit-XL 0.8867 0.9502 0.9241 0.9552 0.8120 0.9056 0.9295 0.9720 0.8824 0.9441 FineEdit-Pro 0.9539 0.9821 0.9521 0.9710 0.8521 0.9185 0.9538 0.9836 0.9245 0.9628
+Table 1: Comparison of LLMs on BLEU and ROUGE-L for LaTeX, DSL, Wiki, Code. Overall data displays average performance among all data categories. The best results are highlighted in bold.
+Method Model Open-Source LaTeX DSL Wiki Code Overall BLEU RG-L BLEU RG-L BLEU RG-L BLEU RG-L BLEU RG-L Zero-shot Gemini 1.5 Flash 0.1745 0.3067 0.9643 0.9787 0.7785 0.8908 0.3882 0.5291 0.5764 0.6763 Gemini 2.0 Flash 0.1304 0.2435 0.4243 0.4328 0.8624 0.9145 0.1426 0.2454 0.3899 0.4591 Llama-3.2-1B 0.3168 0.4371 0.2593 0.3466 0.2291 0.3584 0.1749 0.3196 0.2450 0.3654 Llama-3.2-3B 0.3101 0.4374 0.3274 0.4355 0.2871 0.4073 0.2988 0.4177 0.3059 0.4245 Qwen2.5-3B-Instr 0.5196 0.6344 0.2083 0.2603 0.2845 0.4261 0.3985 0.5138 0.3527 0.4587 Few-shot Gemini 1.5 Flash (2 shot) 0.4811 0.5423 0.0511 0.1167 0.7511 0.8462 0.2388 0.3430 0.3805 0.4621 Gemini 2.0 Flash (2 shot) 0.9099 0.9247 0.0294 0.0406 0.9272 0.9740 0.4719 0.6266 0.5846 0.6415 FineEdit FineEdit-L 0.6823 0.8531 0.8071 0.8730 0.4938 0.6588 0.6707 0.7773 0.6635 0.7906 FineEdit-XL 0.3230 0.4468 0.8050 0.8798 0.4522 0.6333 0.6806 0.7756 0.5652 0.6839 FineEdit-Pro 0.8461 0.8917 0.8123 0.8902 0.6975 0.8286 0.9499 0.9796 0.8265 0.8975
+Table 2: Multi-turn editing results for LaTeX, DSL, Wiki and Code. Overall data displays average performance among all data categories.
 4.2 Performance of Existing Models
@@
 Among all baselines, Gemini 1.5 Flash achieved the highest overall scores, while Mistral-7B-OpenOrca recorded the lowest BLEU and ROUGE-L values.
-Although model size is often a crucial factor, Gemini 2.0 Flash did not surpass Gemini 1.5 Flash in overall effectiveness.
-For instance, despite having more parameters than LLaMA-3.2-1B, Mistral-7B-OpenOrca underperformed in both metrics, highlighting the significance of model architecture and training methods.
-Moreover, while Gemini 2.0 Flash shows superior semantic understanding in the Wiki category—achieving a BLEU score of 0.9133 and a ROUGE-L score of 0.9429—its overall performance remains below that of its counterpart.
+Although model size is typically a critical factor, Gemini 2.0 Flash did not outperform Gemini 1.5 Flash in terms of overall effectiveness.
+For example, despite having more parameters than LLaMA-3.2-1B, Mistral-7B-OpenOrca underperformed on both metrics, highlighting the importance of model architecture and training strategies.
+Additionally, Gemini 2.0 Flash demonstrated superior semantic understanding in the Wiki category, with a BLEU score of 0.9133 and a ROUGE-L score of 0.9429, yet its overall performance remained inferior to that of Gemini 1.5 Flash.
 FineEdit, and in particular its FineEdit-Pro variant, further outperforms all zero-shot baselines.
@@
 We further evaluated few-shot learning on the Gemini models.
-Although few-shot performance notably improved in some categories—for example, in the LaTeX domain, where Gemini 2.0 Flash exhibited a 20% higher BLEU score than in the zero-shot setting—the overall few-shot results still lag behind FineEdit.
-In certain cases, such as the SQL category, few-shot learning made little difference, with BLEU and ROUGE-L scores of only 0.1600 and 0.1814, respectively.
-These findings reinforce the value of our curated benchmark in driving improvements in editing tasks.
-Key Findings: FindEdit demonstrates robust overall effectiveness across Wiki, Code, DSL, and LaTeX categories.
-These results not only position FineEdit as a competitive method for structured editing tasks but also provide valuable insights into how targeted training strategies can elevate model performance in diverse application scenarios.
-Method Model Open-Source LaTeX DSL Wiki Code Overall BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L Zero-shot Gemini 1.5 Flash 0.8665 0.9150 0.8297 0.8555 0.7626 0.8361 0.8551 0.9073 0.8285 0.8819 Gemini 2.0 Flash 0.7413 0.7951 0.4706 0.4964 0.9133 0.9429 0.1339 0.2737 0.5853 0.6519 Llama-3.2-1B 0.5088 0.6108 0.5564 0.6596 0.4413 0.5766 0.4742 0.6072 0.4867 0.6069 Llama-3.2-3B 0.5969 0.6925 0.5747 0.6821 0.5061 0.6384 0.6638 0.7727 0.5862 0.6976 Qwen2.5-3B-Instruct 0.5467 0.6712 0.4107 0.4991 0.4170 0.5699 0.3967 0.5390 0.4492 0.5816 Mistral-7B-OpenOrca 0.3782 0.5770 0.0361 0.1638 0.3608 0.5840 0.3763 0.6447 0.3246 0.5395 Few-shot Gemini 1.5 Flash(2-shot) 0.8742 0.9324 0.0908 0.1190 0.8657 0.9139 0.7412 0.8302 0.7249 0.7845 Gemini 2.0 Flash(2-shot) 0.9464 0.9723 0.1600 0.1814 0.9380 0.9665 0.8327 0.8698 0.8011 0.8302 FineEdit FineEdit-L 0.9311 0.9697 0.9334 0.9615 0.8077 0.9036 0.9296 0.9725 0.8957 0.9504 FineEdit-XL 0.8867 0.9502 0.9241 0.9552 0.8120 0.9056 0.9295 0.9720 0.8824 0.9441 FineEdit-Pro 0.9539 0.9821 0.9521 0.9710 0.8521 0.9185 0.9538 0.9836 0.9245 0.9628
-Table 2: Comparison of LLMs on BLEU and ROUGE-L for LaTeX, DSL, Wiki, Code. Overall data displays average performance among all data categories. The best results are highlighted in bold.
+Although few-shot prompting notably improved performance in some categories, such as the LaTeX domain where Gemini 2.0 Flash achieved a 20% higher BLEU score compared to the zero-shot setting, the overall few-shot results still remained inferior to FineEdit.
+In certain cases, such as the SQL category, few-shot learning provided minimal improvement, achieving BLEU and ROUGE-L scores of only 0.1600 and 0.1814, respectively.
+These findings highlight the effectiveness and importance of our curated benchmark in driving advancements in editing tasks.
 4.3 FineEdit: Supervised Finetuning
@@
 We also compared our models with Gemini’s few-shot prompting approach in real-world scenarios.
-Although in-context learning (ICL) boosts Gemini’s performance in some cases—such as a 8% higher BLEU score in Wiki dataset for Gemini 2.0 Flash—the overall results still lag behind FineEdit-Pro.
-This evidence confirms that our tailored supervised fine-tuning approach yields a more robust and generalizable solution for structured editing tasks.
-Key Findings: FineEdit’s supervised fine-tuning markedly enhances performance.
-FineEdit-L surpasses zero-shot baselines and FineEdit-XL offers comparable gains, while FineEdit-Pro (built on Qwen2.5-3B-Instruct) achieves the highest scores.
-This highlights that robust instruction tuning is more effective than merely scaling model size.
+Although in-context learning (ICL) improves Gemini’s performance in certain cases, such as an 8% increase in BLEU score on the Wiki dataset for Gemini 2.0 Flash, the overall performance remains inferior to FineEdit-Pro.
+This superior performance highlights the effectiveness of our high-quality, rigorously validated InstrEditBench dataset in enabling more robust and generalizable solutions for editing tasks.
+Edit Request 1: Change the brackets in the code to semicolons.
+Original: def test(options): options.data = [] ; def test2(options): options.data = []
+Gemini () def test(options): options.data = None ; def test2(options): options.data = None
+FindEdit Pro () def test(options): options.data = ; ; def test2(options): options.data = ;
+[…]
+[…]
+[…]
+[…]
+Edit Request 3: Replace “Falcon” with “Captain America”.
+Original: In “Captain America: Brave New World,” Sam Wilson, formerly the Falcon, assumes …
+Gemini () In “Captain America: Brave New World,” Sam Wilson, formerly known as the Falcon, assumes …
+FindEdit Pro () In “Captain America: Brave New World,” Sam Wilson, formerly the Captain America, assumes …
+Edit Request 4: Add column created_at TIMESTAMP DEFAULT CURRENT_TIMESTAMP.
+Original: CREATE TABLE worker_salaries (employee_id INT, ...)
+Gemini () ALTER TABLE community_gardens ADD COLUMN created_at TIMESTAMP DEFAULT CURRENT_TIMESTAMP
+FindEdit Pro () CREATE TABLE community_gardens (..., created_at TIMESTAMP DEFAULT CURRENT_TIMESTAMP)
+Table 3: Colour-coded comparison of Gemini and FindEdit Pro responses to four edit requests. ( = incorrect) stands for incorrect editing, while ( = correct) stands for correct editing.
+4.4 Multi-turn Editing Evaluation
+We also evaluated the extended benchmark on FineEdit in the multi-turn setting.
+In this extension, each data instance contains multiple editing requests.
+To simulate real multi-turn scenario, we apply these instructions iteratively: each request is executed on the output produced by the previous one, ensuring that the edits are applied in a cumulative manner.
+We then assess the final output to determine whether it accurately reflects the cumulative effect of all editing instructions after the full sequence of modifications has been applied.
+To assess the performance of multi-turn, we randomly sample 100 multi-turn data for each category, and test them on different models.
+Results show that multi-turn editing leads to consistent performance drops across all domains.
+Specifically, BLEU scores for LaTeX drop from 0.9539 to 0.8461, DSL from 0.9521 to 0.8123, Wiki from 0.8521 to 0.6975, and Code from 0.9538 to 0.9499.
+These results indicate that multi-turn scenarios are substantially more challenging, especially for Wiki and DSL, while code exhibit strong robustness under multi-turn edits.
+This indicates that the decline is the accumulation of errors across turns.
+Each instruction is applied to the output of the previous one, which may already contain small deviations.
+These deviations can propagate through subsequent steps and lead to compounded errors in the final output.
+Despite the degradation in multi-turn settings, FineEdit-Pro achieves an average BLEU of 0.8265, substantially higher than Gemini 1.5 Flash (0.5764) and Gemini 2.0 Flash (0.3899).
+This further demonstrates the effectiveness of our dataset design and its extensibility to diverse editing scenarios.
 4.5 Qualitative Study
-To qualitatively assess the performance of FindEdit, we conduct several studies as shown in Figure 2.
-This figure illustrates eight examples of how FineEdit-Pro and Gemini respond to diverse editing requests.
-In several cases, FineEdit-Pro accurately applies changes—such as adding new columns in DSL or adjusting environment commands—while Gemini often restates the instruction without implementing the intended modifications.
+To qualitatively assess the performance of FindEdit on single-turn editing tasks, we conduct several studies as shown in Table 3.
+This table illustrates eight examples of how FineEdit-Pro and Gemini respond to diverse editing requests.
+In several cases, FineEdit-Pro accurately applies the required changes.
+Specifically, it could correctly add new columns in DSL or adjust environment commands.
+However, Gemini often restates the instruction without actually implementing the intended modifications.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Edit Request 3: Remove the ’TODO’ from the ’TODO’ line.
+[…]
+[…]
+[…]
+Table 4: Color-coded comparison of Gemini and the FineEdit Pro for a multi-turn task with three edit requests.
+( = incorrect) indicates an unsatisfied edit, while ( = correct) indicates a satisfied edit.
 Specifically, both Gemini 1.5 Flash and 2.0 Flash perform well on LaTeX and Wiki tasks, yet they struggle with DSL and Code tasks.
-For example, as shown in Figure 2, FineEdit-Pro correctly identifies the target table and appends a new column named created_at with the data type DEFAULT CURRENT_TIMESTAMP.
+For example, as shown in Table 3, FineEdit-Pro correctly identifies the target table and appends a new column named created_at with the data type DEFAULT CURRENT_TIMESTAMP.
 In contrast, Gemini misinterprets the instruction, merely repeating the edit request rather than applying the intended change.
@@
 Nonetheless, FineEdit is not without shortcomings.
-In the LaTeX example depicted in Figure 2, Gemini accurately locates the subsection{Strengths} and updates it as specified.
+In the LaTeX example depicted in Table 3, Gemini accurately locates the subsection{Strengths} and updates it as specified.
 However, although FineEdit-Pro also identifies and modifies the correct location, it generates the correct response twice, which deviates from the direct editing requirement.
@@
 On the other hand, Gemini occasionally performs better in simpler tasks, such as LaTeX updates.
-Key Findings: FineEdit-Pro demonstrates superior handling of DSL and Code edits compared to Gemini, though minor issues such as duplicate outputs in LaTeX tasks remain.
-Overall, FineEdit’s qualitative performance confirms its robust ability to interpret and execute complex editing instructions.
-Figure 2: Comparison between Gemini and FindEdit Pro response.
+[…]
+[…]
 4.6 Human Evaluation
-To assess whether DiffEval enhances overall dataset quality, we conducted a human evaluation.
-Given that our dataset includes Code and DSL categories—areas closely tied to computer programming—we have three evaluators, each holding at least a Bachelor’s degree in a Computer Science-related field.
+To assess whether DiffEval improves the overall quality of the dataset, we carried out a human evaluation.
+Because the dataset includes Code and DSL categories that require programming expertise, we recruited three evaluators, each with at least a bachelor’s degree in computer science or a related discipline.
 We established the following guidelines to ensure rigorous assessment: (1) Precise Observation: Confirm that the updated content exactly corresponds to the segment specified by the edit request.
@@
 Notably, the DSL dataset experienced the greatest improvement, with quality increasing by over 24% compared to data that did not meet DiffEval’s standards.
-Wiki LaTeX DSL Code
+Threshold Wiki LaTeX DSL Code
 G-score $\geq$ 9 97% 93% 90% 97%
 G-score $<$ 9 87% 89% 66% 83%
-Table 3: Sample performance based on the G-Score
-Key Findings: The DiffEval pipeline significantly improves dataset quality, with Wiki and Code categories achieving high precision, and the DSL category showing over 24% enhancement in quality.
+Table 5: Annotation accuracy across content types under different G-score thresholds.
+4.7 Ablation Study on DiffEval Components
+To better understand the contribution of each component in the DiffEval pipeline, we conducted two ablation experiments. Manual annotation followed the protocol described in Section 4.6.
+Git diff effectiveness.
+We evaluated a reduced pipeline in which G-Eval judged the alignment between the instruction and the edited text without access to git diff.
+From its output, we randomly sampled 100 examples whose G-Score was at least nine and annotated them.
+For comparison, we annotated another 100 examples produced by the full DiffEval pipeline, where G-Eval received the git diff instead of the full edited text.
+Including git diff raised the accuracy from 0.85 to 0.94.
+These results indicate that git diff contributes important structural information for identifying precise alignment between the instruction and the edit.
+G-score threshold selection.
+We also examined the effect of the G-score threshold $\alpha$ used in filtering.
+Setting $\alpha=8$ results in many examples where the core instruction is followed, but this will sometimes introduce unintended formatting changes.
+One common issue is the insertion of extra spaces throughout the text.
+For instance, in a deletion instruction targeting a historical phrase, the phrase was correctly removed, but the resulting diff introduced multiple superfluous spaces across the paragraph.
+This violates the requirement to preserve all formatting outside the instructed change.
+Such formatting issues were significantly reduced when the threshold was increased to $\alpha=9$.
 5 Conclusion
-This work addresses the critical gap in LLMs’ ability to perform precise and targeted text modifications.
-We introduce InstrEditBench, a high-quality benchmark with 20,000+ structured editing tasks across Wiki articles, LaTeX documents, code, and database DSLs, enabling rigorous evaluation of direct editing capabilities.
-To further advance LLMs’ editing proficiency, we propose FineEdit, a specialized model trained on this benchmark.
-Extensive evaluations demonstrate that FineEdit outperforms state-of-the-art models, including GPT-4o, Gemini 2.0, and LLaMa-3.2, with up to 10% improvement compared to Gemini in direct editing task performance.
+We introduce InstrEditBench, a benchmark of over 30k editing tasks spanning Wiki, LaTeX, code, and DSL, aimed at precise instruction-based text editing.
+To ensure supervision quality, we develop DiffEval, an automated pipeline combining structural and semantic filters.
+We further validate our benchmark with FineEdit, a model fine-tuned on InstrEditBench, achieving up to 10% gains over leading models.
+Designed for both single-turn and multi-turn editing, our modular benchmark and pipeline enable broad applicability.
+We will publicly release FineEdit to foster further research in instruction-driven and context-aware text editing.
 6 Limitations
@@
 Controlled Context Evaluation.
-Our benchmark focuses on controlled evaluation context, where it does not yet encompass long-context chain-of-thought scenarios, as smaller LLMs are confined by limited context windows, even though such techniques could be effective in proprietary models
-Appendix A Dataset Generation Prompts
+Our benchmark focuses on controlled evaluation contexts, where it does not yet encompass long-context chain-of-thought scenarios, as smaller LLMs are confined by limited context windows, even though such techniques could be effective in proprietary models.
+Appendix A Additional Implementation Details
+For existing models, we strictly adhere to configurations from their original papers.
+To manage fixed maximum token lengths $L$, if the combined $T_{\text{orig}}$ and $I_{\text{edit}}$ exceed $L$, we partition $T_{\text{orig}}$ into chunks of size $\leq L$, process each chunk independently with the same edit instruction, and concatenate the outputs to form the complete edited text.
+We fine-tune models using Low-Rank Adaptation (LoRA) Hu et al.
+(2021) with $r=8$, $\alpha=32$, and a dropout rate of 0.05, employing the AdamW optimizer with a learning rate of $2\times 10^{-5}$, training for 2 epochs, an effective batch size of 1, and 4 gradient accumulation steps.
+Chunking long context: Many large language models impose a fixed maximum token length $L$ on their input (and sometimes output) sequences.
+Consequently, if the combination of $T_{\text{orig}}$ and $I_{\text{edit}}$ exceeds this limit, we divide the $T_{\text{orig}}$ into smaller chunks of size $\leq L$.
+Each chunk is then processed independently—paired with the same edit request and later concatenated to form the complete edited text.
+This approach ensures that every chunk fits within the model’s token budget, preventing overflow and reducing memory usage while preserving the overall structured editing behavior.
+Fine-Tuning Configuration: We use a LoRA rank of r = 8 and LoRA alpha $\alpha=32$, following the original LoRA paper Hu et al.
+(2021).
+This combination (r = 8, $\alpha=32$) and a learning rate of $2\times 10^{-5}$ are widely used in practice, including in the default settings of the HuggingFace PEFT library.
+It produces a scaling factor of $\alpha/r=4$, which balances training stability and memory efficiency, enabling the model to learn meaningful updates without destabilizing training.
+We set lora_dropout = 0.05, a typical value that helps regularize LoRA updates and reduce overfitting.
+Training and Generation Settings are as follows:
+  • Epochs: 2 epochs, which is generally sufficient for convergence in our editing task.
+  • Gradient Accumulation Steps: 4 (necessary due to a small batch size of 1 and GPU constraints).
+  • Max Chunk Tokens: 2048.
+  • Max Length: 4096.
+  • Generation Settings: temperature = 0.2, top-p = 0.95.
+The token constraints ensure no exceeding of the model’s context window and maintain consistent training across models.
+These parameters reduce randomness while keeping the generated text relevant to the task.
+Decoding and Inference: During generation, we set the temperature to 0.2 and used top-p sampling with a probability of 0.95, then merging outputs from all chunks to produce the final edited text.
+The temperature and top-p settings follow previous editing task studies Cassano et al.
+(2024) to ensure minimal changes rather than creative expansions as our editing tasks require precise.
+Appendix B Data Example
+Table 6 presents representative examples from our benchmark, covering four distinct data categories—WikiText, LaTeX, Code, and Database DSL.
+Each example includes the original content, the user-issued edit request, the resulting edited content, the line-level difference, and the associated G-score indicating edit difficulty.
+[…]
+[…]
+This indicates that the duplicate has been successfully removed.
+Appendix C Dataset Generation Prompts
 We use the following prompts for dataset generation on each domain.
@@
      apart from the editing part.
-We use the following prompt for doing G-eval.
-Appendix B Additional Implementation Details
-Chunking long context: Many large language models impose a fixed maximum token length $L$ on their input (and sometimes output) sequences.
-Consequently, if the combination of $T_{\text{orig}}$ and $I_{\text{edit}}$ exceeds this limit, we divide the $T_{\text{orig}}$ into smaller chunks of size $\leq L$.
-Each chunk is then processed independently—paired with the same edit request and later concatenated to form the complete edited text.
-This approach ensures that every chunk fits within the model’s token budget, preventing overflow and reducing memory usage while preserving the overall structured editing behavior.
-Fine-Tuning Strategy: We use Low-Rank Adaptation (LoRA) [21] to efficiently adapt these models to our task, significantly reducing the number of trainable parameters while preserving their expressive power.
-In all LoRA configurations, We set the rank $r=8$ and scaling $\alpha=32$, and use a dropout probability of 0.05.
-For both Llama-based and Qwen-based models, we apply LoRA to the attention’s projection layers through trainable low-rank matrices.
-We used the AdamW optimizer with a learning rate of $2\times 10^{-5}$, training for 2 epochs, and set the effective batch size of 1 with gradient accumulation steps of 4 due to device limits.
-This strategy not only reduces computational overhead but also enables rapid convergence on our structured editing tasks.
-Preliminary experiments guided the choice of hyperparameters across all three model variants.
-Decoding and Inference: During generation, we set the temperature to 0.7 and used top-p sampling with a probability of 0.9 to balance diversity and coherence.
-Greedy decoding is applied by default if without sampling setting.
-The final edited text is obtained by merging the edited outputs from all chunks.
+We use the following prompts for G-Eval.
+I will give you the content and then the editing request.
+Please Edit the content based on the editing request.
+While Editing, don’t add other words like modified or something.
+Just Edit directly.
+Content: {original_context}
+Editing Request: {edit_request}
+Please return the complete content after editing.
+Don’t skip the empty line and keep the original apart from the editing part.
+Data Category Orignal Content Edit Request Edited Content Difference G-score
+WikiText …As with previous <unk> Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a military unit… Replace “\<unk>\” with “Valkyria” where it appears in the text.
+…As with previous Valkyria Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a military unit… Line 2 differs: Differences: …As with previous [<un -> Val]k[> -> yria] Chronicles games, Valkyria Chronicles III is a tactical role @-@ playing game where players take control of a mili- tary unit… 9
+[…]
+[…]
+[…]
+Table 6: Data examples of different data categories with all attributes (content, edit request, edited content, difference, and G-score).